# DEFormer: DCT-driven Enhancement Transformer for Low-light Image and Dark Vision

### v2

@@ Abstract
Low-light image enhancement restores colors and details of single image and improves high-level visual tasks. However, restoring the lost details in the dark area is a challenge by only relying on the RGB domain. In this paper, we introduce frequency as a new clue into the network and propose a DCT-driven enhancement transformer (DEFormer) framework. First, we propose a learnable frequency branch (LFB) for frequency enhancement contains DCT processing and curvature-based frequency enhancement (CFE) to represent frequency features. In addition, we propose a cross domain fusion (CDF) for reducing the differences between the RGB domain and the frequency domain. Our DEFormer has achieved advanced results in both the LOL and MIT-Adobe FiveK datasets and improved the performance of dark detection.
@@ I Introduction
Low-light images have many dark areas, which result in noise, lose details and bring a negative visual experience. Low-lighe sences make the performance of high-level visual tasks decline (such as object detection[1, 2, 3], face detection[4, 5, 6]). Many low-light image enhancement (LLIE) methods have been proposed recently. Based on the Retinex theory[7], some methods [8, 9, 10] decompose the image into reflection components and illumination components and enhance them. However, traditional methods with manual feature learning has large limitations and poor generalization. Currently, there are also some CNN based [11, 12, 13] and transformer based[14, 15, 16] works proposed for bad scenes such as night, which improve downstream tasks in dark condition.
@@
Figure 1: (a) represents the input, (b) represents the frequency spectrum after DCT processing. The DCT coefficient matrix concentrate the energy of the image signal, so as to realize the effective analysis of the image signal. (c) represents the performance comparison between the SOTA methods in the MIT-Adobe FiveK dataset, the x-axis represents Flops and the y-axis represents SSIM.
Discrete cosine transform (DCT) is an important part of frequency signal analysis, which has been proven effective in computer vision[17, 18, 19]. Inspired by the frequency[20, 21, 22], in this paper we propose a DCT-driven enhancement transformer (DEFormer) for low-light enhancement. We propose a learnable frequency branch (LFB) that contains DCT processing and curvature-based frequency enhancement (CFE) to embed information of DCT. In addition, we propose a cross domain fusion (CDF) for reducing the differences between different domains. We adopt LOL[23] and MIT-Adobe FiveK[24] datasets for low-light enhancement experiments and the results show that DEFormer achieves advanced performance comparing with other state-of-the-art methods. We achieve effective improvements on the ExDark[3] datasets through end-to-end finetuning. The performance comparison of different SOTA methods is shown in Fig. 1 (c).
@@
  • Jointing DCT with transformer, we propose a DCT-driven enhancement transformer (DEFormer) for low-light enhancement and we design a learnable frequency branch (LFB) to embed frequency clue.
  • We propose curvature-based frequency enhancement (CFE) to adaptively focus on frequency-band channels with rich textures. In addition, we propose a cross domain fusion (CDF) to fuse the features of RGB domain and frequency domain.
  • DEFormer has obtained advanced results on public datasets, and end-to-end training of DEFormer and detectors has a outstanding performance improvement for downstream dark detection tasks.
Figure 2: Overview of DEFormer. Two transformer blocks are used to extract shallow features. In the learnable frequency branch (LFB), we introduce frequency clues through DCT processing and curve-based frequency enhancement. In cross domain fusion, the difference between RGB domain and frequency domain is reduced through cross fusion.
@@ II Method
The overview of DEFormer is shown in Fig. 2. Taking a low-light image $I\in R^{H\times W\times 3}$ as input, first we extract the shallow features $F_{0}\in R^{H\times W\times 16}$ through a $3\times 3$ convolution block. Then we apply two transformer blocks to predict a multiple map and an add map respectively and get the feature of the RGB domain $F_{rgb}\in R^{H\times W\times 16}$. In the learnable frequency branch (LFB), we first perform DCT processing on $I$, then divide it into high and low parts with the curvature-based frequency enhancement (CFE) and enhance them respectively, throughout tensor concat we obtain the features $F_{1}\in R^{H/8\times W/8\times 192}$. We reshape $F_{1}$ to $R^{H\times W\times 3}$ and extract features of the frequency domain $F_{f}\in R^{H\times W\times 16}$. We reduce the differences of domain between $F_{rgb}$ and $F_{f}$ by cross fusion in cross domain fusion (CDF). After fusion we apply a transformer block to extract deep features $F_{2}\in R^{H\times W\times 16}$. Finally, we use a $3\times 3$ convolution and Leaky ReLU function to obtain the enhanced image $\hat{I}\in R^{H\times W\times 3}$.
Figure 3: Details of cross domain fusion (CDF). We first complement the information between different domains through cross-fusion, and control the spatial information through a soft attention to reduce noise propagation.
@@ II-B Cross Domain Fusion
Figure 4: Visualization of dark object detection on ExDark dataset. MBLLEN directly predict the enhanced image and apply detector training, and DEFormer uses end-to-end training.
@@ III Experiments
Figure 5: Visualization of different low-light image enhancement methods on LOL dataset. Each row is a different image comparison. We advise that zoom in to observe the details.
@@ III-A Comparison of Low-light Enhancement
We quantitatively compared DEFormer with other state-of-art methods such as RetinexNet [23], Uformer [29], LLFormer [31], and IAT [28]. The results are shown in Table I. Our DEFormer has achieved advanced results in both the LOL and MIT-Adobe FiveK datasets, with only 17.96G of parameters. We visualize the results of some SOTA models and DEFormer in the LOL dataset, as shown in Fig. 5, in contrast our method has high quality in both texture and color.
In addition, we fine-tune DEFormer and YOLOv3 detectors through end-to-end training and we visualize in ExDark datasets, as shown in Fig. 4. Our DEFormer has achieved effective improvement in the ExDark dataset, which is 2.1% higher than baseline in mAP, this shows that our model improves downstream visual tasks.
@@ III-B Ablation Study
We conducted ablation studies on the LOL dataset, and the results are shown in Table II. We directly added the frequency domain information to the network, PSNR and SSIM increased by 0.45 and 0.004 respectively. After adopting the curvature-based frequency enhancement, PSNR and SSIM are 1.02 and 0.012 higher than the baseline, respectively. After using cross domain fusion, PSNR and SSIM have increased by 2.61 and 0.028, respectively. Overall, this ablation study shows that using frequency information as additional clue in the network is an effective solution.
@@ IV CONCLUSIONS
In this paper, we propose DEFormer using frequency as a new clue, which mainly includes learnable frequency branch (LFB) and cross domain fusion (CDF). Extensive experiments show that DEFormer outperforms state-of-the-art methods in low-light enhancement and improves downstream detection tasks well as preprocessing.

### v3

@@ Abstract
Low-light image enhancement restores the colors and details of a single image and improves high-level visual tasks. However, restoring the lost details in the dark area is still a challenge relying only on the RGB domain. In this paper, we delve into frequency as a new clue into the model and propose a DCT-driven enhancement transformer (DEFormer) framework. First, we propose a learnable frequency branch (LFB) for frequency enhancement contains DCT processing and curvature-based frequency enhancement (CFE) to represent frequency features. Additionally, we propose a cross domain fusion (CDF) to reduce the differences between the RGB domain and the frequency domain. Our DEFormer has achieved superior results on the LOL and MIT-Adobe FiveK datasets, improving the dark detection performance.
@@ I Introduction
Low-light images have many dark areas, resulting in noise, loss details and bring a negative visual experience. Low-light sence impaires the performance of high-level visual tasks decline (such as object detection[1, 2, 3], face detection[4, 5, 6] and video task [7, 8]). Many low-light image enhancement (LLIE) methods have recently been proposed. Based on Retinex theory[9], several methods [10, 11, 12] decompose the image into reflection and illumination components and enhance them. However, traditional methods with manual feature learning have large limitations and poor generalization. Recently, several CNN-based [13, 14, 15] and Transformer-based[16, 17, 18] methods have been developed to address challenges in poor lighting conditions, improving dark visual tasks.
@@
Figure 1: (a) represents the input, (b) represents the frequency spectrum with DCT processing. The DCT coefficient matrix concentrate the energy of the image signal. (c) represents the performance comparison between other SOTA methods on the MIT-Adobe FiveK dataset, the x-axis represents Flops and the y-axis represents SSIM.
Discrete cosine transform (DCT) is an significant part of frequency signal analysis, which has been proven effective in computer vision[19, 20, 21]. Inspired by the frequency[22, 23, 24], in this paper we propose a DCT-driven enhancement transformer (DEFormer) for low-light enhancement. We propose a learnable frequency branch (LFB) that contains DCT processing and curvature-based frequency enhancement (CFE) to embed information of DCT. Additionally, we propose a cross domain fusion (CDF) to reduce the differences between different domains. Extensive qualitative and quantitative experiments conducted on the LOL[25] and MIT-Adobe FiveK[26] datasets demonstrate that DEFormer achieves superior performance compared to other state-of-the-art methods. We achieve effective improvements on the ExDark[3] datasets. The performance comparison of different SOTA methods is shown in Fig. 1 (c).
@@
  • Exploring the integration of DCT and Transformers, we propose a DCT-driven enhancement transformer (DEFormer) for low-light enhancement and design a learnable frequency branch (LFB) to incorporate frequency clue.
  • We propose curvature-based frequency enhancement (CFE) to adaptively focus on frequency-band channels with rich textures. Additionally, we propose a cross domain fusion (CDF) technique to integrate features from both the RGB and frequency domains.
  • DEFormer achieved superior results on public datasets, and its end-to-end training with detectors significantly enhances performance in dark detection tasks.
Figure 2: Overview of DEFormer. In the learnable frequency branch (LFB), we introduce frequency clues through DCT processing and curve-based frequency enhancement. In the cross domain fusion (CDF), the difference between the different domains is reduced through cross fusion.
@@ II Method
The overview of DEFormer is shown in Fig. 2. Giving a low-light image $I\in R^{H\times W\times 3}$ as input, we firstly extract the shallow features $F_{0}\in R^{H\times W\times 16}$ through a $3\times 3$ convolution layer. Then, we apply two transformer blocks to predict a multiple map and an add map, respectively, thereby obtaining the features of the RGB domain $F_{rgb}\in R^{H\times W\times 16}$. In the learnable frequency branch (LFB), we perform DCT processing on $I$ and divide it into high and low parts using the curvature-based frequency enhancement (CFE), enhancing each part separately, throughout tensor concat we obtain the features $F_{1}\in R^{H/8\times W/8\times 192}$. We reshape $F_{1}$ to $R^{H\times W\times 3}$ and extract features of the frequency domain $F_{f}\in R^{H\times W\times 16}$. We reduce the differences of domain between $F_{rgb}$ and $F_{f}$ in cross domain fusion (CDF). Following the fusion, we apply a transformer block to extract deep features $F_{2}\in R^{H\times W\times 16}$. Finally, we use a $3\times 3$ convolution and Leaky ReLU function to obtain the enhanced image $\hat{I}\in R^{H\times W\times 3}$.
Figure 3: Details of cross domain fusion (CDF). We first complement the information between different domains through cross fusion, controling the spatial information through a soft attention to reduce noise propagation.
@@ II-B Cross Domain Fusion
Figure 4: Visualization of dark object detection on the ExDark dataset. MBLLEN directly predict the enhanced image and apply detector training, DEFormer uses end-to-end training.
@@ III Experiments
Figure 5: Visualization of different low-light image enhancement methods on the LOL dataset. Each row is a different image sample. We advise that zoom in to observe the details.
@@ III-A Comparison of Low-light Enhancement
We quantitatively compared DEFormer with other state-of-art methods such as RetinexNet [25], Uformer [31], LLFormer [33] and IAT [30], the results are shown in Table I. Our DEFormer has achieved superior results on the LOL and MIT-Adobe FiveK datasets, with only 17.96G on the FLOPs. We visualized the results of several SOTA models and DEFormer on the LOL dataset, as shown in Fig. 5, in contrast our method had high quality in both texture and color.
In addition, we fine-tuned the DEFormer and YOLOv3 detectors through end-to-end training and visualized on the ExDark datasets, as shown in Fig. 4. Our DEFormer has achieved an effective improvement on the ExDark dataset, which is 2.1% higher than the baseline in mAP. This demonstrated that our model improves downstream visual tasks.
@@ III-B Ablation Study
We conducted ablation studies on the LOL dataset, the results are shown in Table II. We directly added the frequency domain information to the network, PSNR and SSIM increased by 0.45 and 0.004 respectively. After adopting the curvature-based frequency enhancement, PSNR and SSIM are 1.02 and 0.012 higher than the baseline, respectively. After using cross domain fusion, PSNR and SSIM have increased by 2.61 and 0.028 respectively. Overall, this ablation study demonstrated that using frequency information as additional clue in the network is an effective solution.
@@ IV CONCLUSIONS
In this paper, we propose DEFormer using frequency as a new clue, which mainly includes learnable frequency branch (LFB) and cross domain fusion (CDF). Extensive experiments demonstrate that DEFormer outperforms state-of-the-art methods in low-light enhancement and improves downstream detection tasks well as preprocessing.
V Acknowledgments
This work was supported by Special Project of the National Natural Science Foundation of China (62441614), Anhui Province Key R&D Program (202304a05020068) and General Programmer of the National Natural Science Foundation of China (62376084).